OpenStudy (anonymous):

p,q,r,s are real numbers.show p^4 + q^4 >= 2(p^2)(q^2). Using that show p^4 + q^4 + r^4 + s^4 >= 4pqrs

ganeshie8 (ganeshie8):

use AM-GM inequality

OpenStudy (anonymous):

first expression can be shown but second one cannot

ganeshie8 (ganeshie8):

\[\large \dfrac{p^4+q^4}{2} \ge \sqrt{p^4q^4}\]

ganeshie8 (ganeshie8):

\[\large \dfrac{p^4+q^4}{2} \ge p^2q^2\]

ganeshie8 (ganeshie8):

similarly \[\large \dfrac{r^4+s^4}{2} \ge r^2s^2\]

OpenStudy (anonymous):

right i did that.then what? I added them but how to get 4pqrs?

ganeshie8 (ganeshie8):

Adding them gives you : \[\large \dfrac{p^4 + q^4}{2} + \dfrac{r^4+s^4}{2} \ge p^2q^2 + r^2s^2\]

ganeshie8 (ganeshie8):

\[\large p^4+q^4+r^4+s^4 \ge 4\left( \dfrac{p^2q^2 + r^2s^2}{2} \right)\] \[\large p^4+q^4+r^4+s^4 \ge 4\left( \sqrt{p^2q^2*r^2s^2} \right)\]

OpenStudy (anonymous):

oh right beautiful thanks again ganeshie8

jagr2713 (jagr2713):

nice @ganeshie8